Daily Summary – May 21st

Crew 181 Daily Summary Report

May 21st, 2017


MDRS Daily Summary Report for SOL 8


Summary Title: A Rainy Day on Mars

Mission Status: A day full of Settlers of Catan and getting too excited about a potential thunderstorm.

Sol Activity Summary: Janet and Charlie had another late night troubleshooting the observatory. Luckily, they are starting to make progress! In the early AM, Janet, Charlie, and Avishek went to Gallieo Rd (see EVA report for more).

Around 2-3pm, we endured our first Martian rain storm that brought heavy winds. All crewmembers took advantage of the tunnel system to get a sense of the storm and make sure there wasn’t too much damage afflicted to the Hab and the buildings. We then had tuna melt quesadillas and played Settlers of Catan while the storm raged on.

After a few rounds of Catan, Charlie had remained undefeated for the day and nobody wanted to play a third round. We checked our internet and realized our usage was up (still unclear what caused this). We are now writing our reports that will be sent in the early AM tomorrow.

Look Ahead Plan: More observatory tonight if it’s clear enough. We will be out of sim tomorrow (aside from a morning EVA) as we prepare for our Martian invaders, Ozzy and Jack! We will not have reports tomorrow, as it will be a tiring day for all involved. Avishek also leaves us tomorrow, which we are all sad to lose a crewmember mid way through our sim (but PhD work can’t wait, apparently.)

Anomalies In Work: Water heater continues to not work (more info in the engineering report). We also lost Internet connection around noon and continue to not have it. We are unsure if this is because of an earlier service manual for the water heater that Charlie downloaded, or if someone accidentally had something connected to wifi. There does not appear to be any issue with the router, but we have tried all options with it. During the rainstorm, we realized a panel of the Hab seems to be flapping when winds from the north gust against it, creating a loud heavy sound.

Weather: Partly cloudy all day leading up a windy and rainy storm around 2-3pm.

Crew Physical Status: Good spirits all around, a little sad that Avishek will be leaving us tomorrow.

EVA: Morning EVA went to Galileo Rd. Afternoon EVA was canceled.

Reports to be Filed: journalist report, engineering report, EVA report, and EVA request.

Support Requested: Unless you can send out a Martian technician to fix or replace the water heater, we do not have any requests.


Prepared by Cassandra



Cassandra Klos

Commander, Crew 181